OpenStudy (anonymous):

What is the vertex for the graph of y = 2x^2 + 12x + 10? (−3, −8) (−3, −44) (3, 64) (3, 58)

OpenStudy (helder_edwin):

\[ \large y=2x^2+12x+10 \] \[ \large y-10+18=2(x^2+6x+9) \] \[ \large y+8=2(x+3)^2 \]

OpenStudy (helder_edwin):

so the vertex is (-3,-8)

OpenStudy (anonymous):

thank you :)

OpenStudy (helder_edwin):

u r welcome did u understand everything?